Question 5 What is the pH of a 0.375 M solution of benzoic acid? Ka=6.3E-5 11.7 0.43 2.3 5.1
Answer #1

Reaction is--C6H5COOH(aq) = C6H5COO-(aq) + H+(aq)

Initial: 0.375 0 0

Change: -x +x +x

Equilibrium: 0.375-x x x

Ka = [C6H5COO-][H+]/[C6H5COOH]

=> x^2/(0.375-x) = 6.3*10^-5

=> x = 0.004829 M

So, [H+] = x = 0.004829 M

pH = -log[H+] = -log(0.004829) = 2.3

If you have any questions please comment

If you satisfied with the solution please rate it thanks...